OpenStudy (goldrush18):

For the function f:x maps to 4x-x^3 determine: a. the coordinates of the stationary points b. the nature of each of the stationary points

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

yep!

OpenStudy (anonymous):

f(x)=4x-x^3 dy/dx=4-3x^2

OpenStudy (anonymous):

solve for when dy/dx=0

OpenStudy (goldrush18):

from there what do i do because i can't get the coordinates

OpenStudy (anonymous):

hence, x^2=SR4/3

OpenStudy (anonymous):

when y=0 of course

OpenStudy (anonymous):

well, y intercept (c) is 4 right?

OpenStudy (anonymous):

so co-ords are (0,4)

OpenStudy (anonymous):

that's the maxima, stationary

OpenStudy (anonymous):

funny thing is, i'm revising for this topic too

OpenStudy (anonymous):

when you asked the question lol

OpenStudy (anonymous):

where are you from?

OpenStudy (goldrush18):

cool but i still stuck here at the point when i get dy/dx=4-3^2

OpenStudy (goldrush18):

im from Jamaica u?

OpenStudy (anonymous):

ooh cool! I live in Hong Kong

OpenStudy (goldrush18):

how do u know that dy/dx=0

OpenStudy (anonymous):

\[f(x)=4x-x^3\] \[dy/dx=-3x^2+4=0\]

OpenStudy (anonymous):

what's dy/dx? it's the gradient right? so when it's at maxima or minima, gradient is always 0

OpenStudy (anonymous):

and since this is a negative quadratic graph, it's maxima

OpenStudy (anonymous):